[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Thu Jul 27, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Thu Jul 27 20:25:45 UTC 2017


HSU Shift Log for Thu Jul 27 2017 16:25:37 GMT-0400 (EDT)

Shift: H

Forecaster: D Zelinsky

Technical Issues Summary:

ATCF issues: TSB working on NETAPP migration. Originally planned to migrate  
ATCF mount points today, but this was delayed. ATCFSVR2 (development sever)  
migrated today.

N-AWIPS issues: None

AWIPS-2/NCP issues: None

Other Techincal issues: None

Miscellaneous issues: Repeated request for 5-min full disk tomorrow.  
Request for today was overruled by a higher priority. Also added a request  
for meso area beginning at 11Z tomorrow for Hilary. Communicated to SDM  
that preference is for 5 minute full disk.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170727/9f2f6462/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list